OpenStudy (anonymous):

In triangle ABC, Angle B measure 5 degrees greater than angle A, angle C measures 7 degrees greather than angle B. Find the measure of of the three angles PLZZZ HELPP !!!!

OpenStudy (amilapsn):

Hint: A+B+C=180\(^o\)

OpenStudy (anonymous):

Let Angle B = A + 5; Let Angle C = B + 7 or (A + 5) +7....find one angle first then use it to figure out the other two angles. check your work afterwards to see if it equals 180.

OpenStudy (anonymous):

How do I find the measure of them ? @mathconceptz